2. Use the method for solving homogeneous equations to solve the following differential equation 8(x2 + y2)dx + 9xydy = 0
Answer #1

(2 Here the differential equation 8 (n2 + y2) dnt any dy = 0 degree are game. So it is homogeneous equation 8 (n2 + y2)dn t aqun du -8-8v2 -912 an D dn n audr an 18 +1402) n 9 v dv 8+1712 taking both side integration we go qvar dn 8+1712 n. For left기고 음 D 려 C ) 36 ( Since z = 8+1702) 3개의 * 8+1V2= 드 건 V2= 두 - 8 v2 C 177 344 8 17 2 C 8 17 (since v= y) 거2 17 349 8 17 2 + 1 유Check the last box... this is the solution of the given homogeneous equation